# Non-commutative rings with infinitely many maximal subrings

### v1

@@ 2. Preliminaries
It is clear that if $R$ is of the forms that is given in the statement of the lemma, then $R$ is a maximal subring of $T$. Conversely, let $R$ be a maximal subrings of $T$. If $0\times B\subseteq R$, then it is not hard to see that $R=A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$. Similarly if $A\times 0\subseteq R$, then $R=A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$. Hence assume that $R$ is a maximal subring of $T$ which does not contain $A\times 0$ and $0\times B$. Let $M=(A\times 0)\cap R=\{(a,0)\ |\ a\in A_{1}\}=A_{1}\times 0$, where $A_{1}$ is a proper subset of $A$ and $N=(0\times B)\cap R=\{(0,b)\ |\ b\in B_{1}\}=0\times B_{1}$, where $B_{1}$ is a proper subset of $B$. Clearly $M$ and $N$ are ideals of $R$ and $M+N=A_{1}\times B_{1}$. Also note that since $A\times 0\nsubseteq R$ and $R$ is a maximal subring of $T$, we have $R+(A\times 0)=T$ and therefore $R/M\cong B$ and similarly $R/N\cong A$. Now we have two cases (note $(1,1)\in R$): $(i)$ $(1,0)\in R$ which equivalent to $(0,1)\in R$. Therefore in this case $M+N=R$ and thus $R=A_{1}\times B_{1}$. Hence $A_{1}$ and $B_{1}$ are proper subrings of $A$ and $B$, respectively. Clearly $R=A_{1}\times B_{1}\subsetneq A\times B_{1}\subsetneq T$, which contradicts to the maximality of $R$. $(ii)$ hence assume that $(1,0)$ and $(0,1)$ are not in $R$. Thus $(1,1)\notin M+N$ and therefore $M+N$ is a proper ideal of $R$ which contains $M$ and $N$. Hence, by $R/M\cong B$ and $R/N\cong A$, we deduce that $A$ and $B$ have epimorphic images that are isomorphic as ring which is impossible. Thus we are done. ∎
@@ 3. Rings with Infinitely Many Maximal Subrings
In this section we study the number of maximal subrings for general non-commutative rings, especially for division rings, simple rings and also $K$-algebras over an infinite field $K$. If $E$ is a field with prime subfield $F$, then by [10, Corollary 1.5], whenever $E$ has only finitely many maximal subrings, then $F\cong\mathbb{Z}_{p}$ for some prime number $p$ and $E$ is algebraic over $F$ (for characterization of such fields see [2]). Moreover, if $E$ is not an absolutely algebraic field, then $|RgMax(E)|\geq max\{tr.deg(E/F),|E|\}$, by [10, Corollary 1.5]. For division rings we have the next result.
@@
Let $D$ be a non-commutative division ring. If $D$ has a maximal subring $R$ which is a division ring, then $D$ has infinitely many maximal subrings. Consequently, if $D$ has only finitely many maximal subrings, then each $x\in D\setminus C_{D}(D)$ is not algebraic over $C_{D}(D)$ (i.e., $C_{D}(D)$ is algebraically closed in $D$) and $C_{D}(R)=C_{D}(D)$ for each maximal subring $R$ of $D$.
@@
Assume that $R$ be a maximal subring of $D$ which is a division ring. Then by [17, Theorem 3], there exist infinitely many subrings of the form $xRx^{-1}$, where $x\in D\setminus\{0\}$. It is evident that for each $0\neq x\in D$, $xRx^{-1}$ is a maximal subring of $D$. Thus $D$ has infinitely many maximal subrings. For the next part, assume that $D$ has only finitely many maximal subrings, then by the first part we conclude that each maximal subring of $D$ is not a division ring. Now if $x$ is a non-centeral element of $D$ which is algebraic over the center of $D$, then by [5, Corollary 2.10], $D$ has a maximal subring which is a division ring which is impossible. Hence the center of $D$ is algebraically closed in $D$. For the final part note that if $R$ is a maximal subring of $D$ and $\alpha\in C_{D}(R)$ is not a central element of $D$, then clearly $R=C_{D}(\alpha)$, i.e., $R$ is a division ring which is impossible too. Thus $C_{D}(R)=C_{D}(D)$. ∎
@@
Let $x\in D\setminus F$ and $[F(x):F]<|F|$. Then similar to the proof of [21, Theorem 4.20], we deduce that $F(x)$ is algebraic over $F$. Hence $x$ is algebraic over $F$ and thus we are done by the previous result. ∎
@@
Let $D$ be an existentially complete non-commutative division ring over a field $K$. Then $D$ has infinitely many maximal subrings.
@@
For more results first we need to characterize the maximal subrings of $R\times R$ for a simple ring $R$, as follow.
@@
Let $R$ be a simple ring. Then $S$ is a maximal subring of $T:=R\times R$ if and only if $S$ satisfies exactly one of the following conditions:
@@
    There exist $\sigma_{1},\sigma_{2}\in Aut(R)$ such that $S=\{(\sigma_{1}(x),\sigma_{2}(x))\ |\ x\in R\}$.
@@
First assume that $S$ is a subring of $T$ which is of the forms given in $(1)$ or $(2)$. It is clear that if $S$ is of the forms given in $(1)$, then $S$ is a maximal subring of $T$. Hence assume that $S$ is of the form in $(2)$. Define $\sigma:R\longrightarrow S$, by $\sigma(x)=(\sigma_{1}(x),\sigma_{2}(x))$, then one can easily check that $\sigma$ is an isomorphism and therefore $S\cong R$ is a simple ring. To show $S$ is a maximal subring of $T$, we first prove that $S[(1,0)]=T$. Clearly $(0,1)\in S[(1,0)]$, for $(1,1)\in S$. Now let $(a,b)\in T=R\times R$, thus there exist $x,y\in R$ such that $\sigma_{1}(x)=a$ and $\sigma_{2}(y)=b$, for $\sigma_{1},\sigma_{2}\in Aut(R)$. Thus by definition of $S$ we conclude that $(\sigma_{1}(x),\sigma_{2}(x))=(a,\sigma_{2}(x))\in S$ and $(\sigma_{1}(y),\sigma_{2}(y))=(\sigma_{1}(y),b)\in S$. Therefore $(a,b)=(a,\sigma_{2}(x))(1,0)+(\sigma_{1}(y),b)(0,1)\in S[(1,0)]$. Thus $S[(1,0)]=T$. Now we prove that for each $(x,y)\in T\setminus S$, we have $S[(x,y)]=T$, which shows that $S$ is a maximal subring of $T$. Since $\sigma_{2}\in Aut(R)$, there exists $c\in R$ such that $\sigma_{2}(c)=y$. Therefore $(\sigma_{1}(c),y)\in S\subseteq S[(x,y)]$ and therefore $(x-\sigma_{1}(c),0)\in S[(x,y)]$. Since $(x,y)\notin S$, we infer that $x_{0}:=x-\sigma_{1}(c)\neq 0$. Thus $Rx_{0}R=R$, for $R$ is a simple ring. Hence there exist $n\in\mathbb{N}$ and $r_{i},s_{i}\in R$, $1\leq i\leq n$, such that $\sum_{i=1}^{n}r_{i}x_{0}s_{i}=1$. From $\sigma_{1}\in Aut(R)$, we deduce that there exist $a_{i},b_{i}\in R$, $1\leq i\leq n$, such that $\sigma_{1}(a_{i})=r_{i}$ and $\sigma_{1}(b_{i})=s_{i}$. Therefore
@@
Thus $T=S[(1,0)]\subseteq S[(x,y)]$ and therefore $S[(x,y)]=T$. Hence $S$ is a maximal subring of $T$.
Conversely, assume that $S$ be a maximal subring of $T$. Suppose $I=R\times 0$ and $J=0\times R$. If either $I\subseteq S$ or $J\subseteq S$, then $S$ is of the forms which is given in $(1)$. Thus assume that $I$ and $J$ are not contained in $S$. Therefore $I_{1}=I\cap S=\{(x,o)\ |\ x\in S_{1}\}=S_{1}\times 0$ and $J_{1}=J\cap S=\{(0,y)\ |\ y\in S_{2}\}=0\times S_{2}$, where $S_{1}$ and $S_{2}$ are proper subsets of $R$. Clearly $I_{1}+J_{1}=S_{1}\times S_{2}\subseteq S$. Also note that since $S$ is a maximal subring of $T$ which does not contain $I$ and $J$, we deduce that $S+I=T=S+J$ and thus $S/I_{1}\cong R\cong S/J_{1}$. Thus $I_{1}$ and $J_{1}$ are maximal ideals of $S$. Now we have two cases: $(a)$ if $(1,0)\in S$ (which is equivalent to $(0,1)\in S$), then $(1,1)=(1,0)+(0,1)\in I_{1}+J_{1}$, which means $I_{1}+J_{1}=S$ and therefore $S=S_{1}\times S_{2}$. Thus $S_{1}$ and $S_{2}$ are proper subring of $R$. But clearly $S\subsetneq R\times S_{2}\neq T$, which contradicts the maximality of $S$. $(b)$ Hence $(1,0),(1,0)\notin S$. Thus $I_{1}+J_{1}=S_{1}\times S_{2}$ is a proper ideal of $S$ which contains $I_{1}$ and $J_{1}$. But as we see earlier, $I_{1}$ and $J_{1}$ are maximal ideals of $S$. Therefore we immediately conclude that $I_{1}=J_{1}$. Thus $I_{1}=J_{1}=0$ and therefore $S\cong R$. Let $\sigma:R\longrightarrow S$ be a ring isomorphism, and for each $x\in R$, let $\sigma(x)=(\sigma_{1}(x),\sigma_{2}(x))$. Then clearly $\sigma_{1}$ and $\sigma_{2}$ are ring endomorphisms of $R$. Since $R$ is a simple ring we conclude that $\sigma_{1}$ and $\sigma_{2}$ are one-one. Finally note that since $S=\{(\sigma_{1}(x),\sigma_{2}(x))\ |\ x\in R\}$ and $(1,0)\notin S$, we deduce that $S[(1,0)]=T$. This immediately implies that $T=S+S(1,0)$ (note, $(1,0)$ is a central idempotent of $T$). To show that $\sigma_{1}$ and $\sigma_{2}$ are onto, let $x\in R$, then $(x,0)\in S+S(1,0)$. Therefore there exist $a,b\in R$ such that $(x,0)=(\sigma_{1}(a),\sigma_{2}(a))+(\sigma_{1}(b),\sigma_{2}(b))(1,0)$, which immediately shows that $\sigma_{1}(b)=x$ and therefore $\sigma_{1}$ is onto. Similarly, $\sigma_{2}$ is onto and hence we are done. ∎
Next we have the following main results in this paper for simple rings.
@@
Let $T$ be a simple ring which is not a division ring. Then $T$ has only finitely many maximal subring if and only if $R$ is finite (i.e., $T\cong\mathbb{M}_{n}(F)$, where $F$ is a finite field and $n\geq 2$)
@@
First we show that $Max_{l}(T)$ is finite. For otherwise let $M_{1},M_{2},\ldots$ be distinct maximal left ideals of $T$. Then for each $i$, by $(1)$ of Theorem 2.7, $A_{i}=\mathbb{I}_{T}(M_{i})$ is a maximal subring of $T$ which contains $M_{i}$ (note, for each $i$, $M_{i}$ is not an ideal of $T$). Since for $i\neq j$, $M_{i}+M_{j}=T$, we immediately conclude that $A_{i}\neq A_{j}$. Thus $T$ has infinitely many maximal subrings which is absurd. Thus $Max_{l}(T)=\{M_{1},\ldots,M_{n}\}$. Now since $J(T)=0$, we deduce that $M_{1}\cap\cdots\cap M_{n}=0$. Thus $T$ can be embedded in $T/M_{1}\times\cdots\times T/M_{n}$. Hence $T$ is a left Artinian ring. Therefore $T\cong\mathbb{M}_{n}(D)$ for some division ring $D$ and $n\geq 2$ (note, $T$ is not a division ring). If $D$ is an infinite division ring, then $\mathbb{M}_{n}(D)$, where $n\geq 2$, has infinitely many maximal left (right) ideals, by Corollary 2.4. Thus $D$ is finite and therefore $D$ is a finite field. Hence $T$ is a finite ring. ∎
We remind the reader that if $K$ is a field, then $K\times K$ has only finitely many maximal subrings if and only if $K$ is finite, see [10, Corollary 3.5]. Now the following generalization is in order.
@@
Let $R$ be a simple ring. Then $R\times R$ has finitely many maximal subrings if and only if $R$ is finite.
@@
It is clear that if $R$ is finite then $R\times R$ has only finitely many maximal subrings. Conversely, assume that $R\times R$ has only finitely many maximal subrings. Thus $R$ has only finitely many maximal subrings. If $R$ is not a division ring, then we are done by the previous theorem. Hence assume that $R$ is a division ring and $R\times R$ has only finitely many maximal subrings. By the preceding comment we may assume that $R$ is not a field and therefore $C_{R}(R)\subsetneq R$. Now suppose that $R$ is infinite. Then we have two cases: $(a)$ If there exists a non-central element $x$ of $R$ which is algebraic over the center of $R$, then by Proposition 3.1, $R$ has infinitely many maximal subrings. Thus $R\times R$ has infinitely many maximal subring too, which is impossible. $(b)$ Hence assume that each non-central element $x$ of $R$ is not algebraic over the center of $R$. For each $n\geq 1$ and $a\in D$, let $\sigma_{n}(a)=x^{n}ax^{-n}$ be the inner automorphism of $R$ induced from $x^{n}$. If $\sigma_{n}=\sigma_{m}$ for $n>m$, then $\sigma_{1}^{n-m}=1$ and therefore $x^{n-m}$ is a central element of $R$. This immediately implies that $x$ is integral over the center of $R$, which contradict to the assumption in this case. Thus $\sigma_{n}\neq\sigma_{m}$ for each $n\neq m$. Finally note that by Theorem 3.4, for each $n\geq 1$, the set $S_{n}=\{(x,\sigma_{n}(x))\ |\ x\in R\}$ is a maximal subring of $R\times R$ and clearly for $n\neq m$, we have $S_{n}\neq S_{m}$. Therefore $R\times R$ has infinitely many maximal subrings, which is absurd. Thus $R$ is a finite and hence we are done. ∎
@@
By [12, Theorem 6.5.11], $D$ has at least $2^{\aleph_{0}}$ automorphisms. Hence by Theorem 3.4 for each $\sigma\in Aut(D)$, $S_{\sigma}:=\{(x,\sigma(x))\ |\ x\in D\}$ is a maximal subring of $D\times D$. It is clear that $S_{\sigma}\neq S_{\sigma^{\prime}}$, whenever $\sigma$ and $\sigma^{\prime}$ are two distinct automorphisms of $D$. Hence $D\times D$ has at least $2^{\aleph_{0}}$ maximal subrings and hence we are done. ∎
@@
It is clear that if $(1)$ holds then $T$ has infinitely many maximal subrings. Now assume that $(2)$ holds, thus $T\cong R_{1}\times R_{1}$ and therefore by Proposition 3.6, we conclude that $T$ has infinitely many maximal subrings. Conversely, assume that $T$ has infinitely many maximal subrings, we prove that $(1)$ or $(2)$ holds. Suppose that $(2)$ does not hold, i.e., $R_{1}$ and $R_{2}$ are not isomorphic rings. Thus $R_{1}$ and $R_{2}$ are homomorphically non-isomorphic rings. Therefore by Lemma 2.9, we infer that each maximal subrings of $T$ is of the form $S_{1}\times R_{2}$ or $R_{1}\times S_{2}$, where $S_{i}$ is a maximal subring of $R_{i}$. Since $T$ has infinitely many maximal subrings, we immediately conclude that either $R_{1}$ or $R_{2}$ has infinitely many maximal subrings and hence we are done. ∎
For a ring $T$, let us denote the sets $Max_{l}(T)\setminus Max(T)$ by $Max_{l}^{\prime}(T)$, similarly $Max_{r}^{\prime}(T):=Max_{r}(T)\setminus Max(T)$ and $Max^{\prime}(T):=Max(T)\setminus Max_{l}(T)=Max(T)\setminus Max_{r}(T)$. Note that if $P$ is a prime ideal of a ring $T$ and $A$ be a left (resp. right) ideal and $J$ be an ideal of $T$ such that $P=I\cap J$, then either $P=I$ or $P=J$. To see this note that clearly $JI\subseteq I\cap J\subseteq P$ (resp. $IJ\subseteq I\cap J\subseteq P$), therefore $I\subseteq P$ or $J\subseteq P$. Hence $P=I$ or $P=J$.
@@
    $Max_{l}^{\prime}(T)$ and $Max_{r}^{\prime}(T)$ are finite sets and in fact $|Max_{l}^{\prime}(T)|\leq|RgMax(T)|$ and $|Max_{r}^{\prime}(T)|\leq|RgMax(T)|$.

### v2

@@ 2. Preliminaries
It is clear that if any subring of the forms described above is maximal in $T$. Conversely, let $R$ be a maximal subrings of $T$. If $0\times B\subseteq R$, then it is not hard to see that $R=A_{1}\times B$, where $A_{1}$ is a maximal subring of $A$. Similarly, if $A\times 0\subseteq R$, then $R=A\times B_{1}$, where $B_{1}$ is a maximal subring of $B$. Hence assume that $R$ contains neither $A\times 0$ nor $0\times B$. Let $M:=(A\times 0)\cap R=\{(a,0)\ |\ a\in A_{1}\}=A_{1}\times 0$, where $A_{1}$ is a proper subset of $A$ and $N:=(0\times B)\cap R=\{(0,b)\ |\ b\in B_{1}\}=0\times B_{1}$, where $B_{1}$ is a proper subset of $B$. Clearly $M$ and $N$ are ideals of $R$ and $M+N=A_{1}\times B_{1}$. Also note that since $A\times 0\nsubseteq R$ and $R$ is a maximal subring of $T$, we have $R+(A\times 0)=T$; Consequently, $R/M\cong B$. Similarly, $R/N\cong A$. Now we distinguish two cases (recall that $(1,1)\in R$): $(\bf i)$ $(1,0)\in R$ (equivalently $(0,1)\in R$). Then $M+N=R$, so $R=A_{1}\times B_{1}$. Hence $A_{1}$ and $B_{1}$ are proper subrings of $A$ and $B$, respectively. But then $R=A_{1}\times B_{1}\subsetneq A\times B_{1}\subsetneq T$, contradicting maximality of $R$. $({\bf ii})$ $(1,0)\notin R$ and $(0,1)\notin R$. Thus $(1,1)\notin M+N$, so $M+N$ is a proper ideal of $R$ contains $M$ and $N$. From $R/M\cong B$ and $R/N\cong A$, we obtain that $A$ and $B$ have isomorphic homomorphic images, which is impossible, because $A$ and $B$ are homomorphically non-isomorphic. Since both cases lead to a contradiction, our assumption that $R$ contains neither $A\times 0$ nor $0\times B$ cannot hold. Thus $R$ must be one of the two forms stated in the lemma. ∎
@@ 3. Rings with Infinitely Many Maximal Subrings
In this section we study the number of maximal subrings for general noncommutative rings, with particular emphasises on division rings, simple rings and $K$-algebras over an infinite field $K$. Let $E$ be a field with prime subfield $F$. By [10, Corollary 1.5], if $E$ has only finitely many maximal subrings, then $F\cong\mathbb{Z}_{p}$ for some prime number $p$ and $E$ is algebraic over $F$ (for a full characterization of such fields see [2]). Moreover, if $E$ is not an absolutely algebraic field, then $|\operatorname{RgMax}(E)|\geq\max\{\operatorname{tr.deg}(E/F),|E|\}$, by [10, Corollary 1.5]. For division rings we obtain the following result.
@@
Let $D$ be a noncommutative division ring. If $D$ has a maximal subring $R$ which is a division ring, then $D$ has infinitely many maximal subrings. Consequently, if $D$ has only finitely many maximal subrings, then each $x\in D\setminus C_{D}(D)$ is not algebraic over $C_{D}(D)$ (i.e., $C_{D}(D)$ is algebraically closed in $D$) and $C_{D}(R)=C_{D}(D)$ for each maximal subring $R$ of $D$.
@@
Assume that $R$ is a maximal subring of $D$ which is a division ring. Then by [17, Theorem 3], there exist infinitely many subrings of the form $xRx^{-1}$, where $x\in D\setminus\{0\}$. It is evident that for each $0\neq x\in D$, $xRx^{-1}$ is a maximal subring of $D$. Thus $D$ has infinitely many maximal subrings. For the next part, assume that $D$ has only finitely many maximal subrings, then by the first part we conclude that each maximal subring of $D$ is not a division ring. Now, if $x$ is a non-centeral element of $D$ which is algebraic over the center of $D$, then by [5, Corollary 2.10], $D$ has a maximal subring which is a division ring which is impossible. Hence the center of $D$ is algebraically closed in $D$. For the final part note that if $R$ is a maximal subring of $D$ and $\alpha\in C_{D}(R)$ is not a central element of $D$, then clearly $R=C_{D}(\alpha)$, i.e., $R$ is a division ring which is absurd. Thus $C_{D}(R)=C_{D}(D)$. ∎
@@
Let $x\in D\setminus F$ and $[F(x):F]<|F|$. Then analogous to the proof of [21, Theorem 4.20], we deduce that $F(x)$ is algebraic over $F$. Hence $x$ is algebraic over $F$ and thus we are done by the previous result. ∎
@@
Let $D$ be an existentially complete noncommutative division ring over a field $K$. Then $D$ has infinitely many maximal subrings.
@@
To obtain further results, we first need to characterize the maximal subrings of $R\times R$ where $R$ is a simple ring. This done as follows.
@@
Let $R$ be a simple ring. Then $S$ is a maximal subring of $T:=R\times R$ if and only if $S$ satisfies exactly in one of the following conditions:
@@
    There exist $\sigma_{1},\sigma_{2}\in\operatorname{Aut}(R)$ such that $S=\{(\sigma_{1}(x),\sigma_{2}(x))\ |\ x\in R\}$.
@@
First assume that $S$ is a subring of $T$ which is of the forms given in $(1)$ or $(2)$. It is clear that if $S$ is of the forms given in $(1)$, then $S$ is a maximal subring of $T$. Hence assume that $S$ is of the form in $(2)$. Define $\sigma:R\longrightarrow S$, by $\sigma(x)=(\sigma_{1}(x),\sigma_{2}(x))$, then one can easily check that $\sigma$ is an isomorphism and therefore $S\cong R$ is a simple ring. To show $S$ is a maximal subring of $T$, we first prove that $S[(1,0)]=T$. Clearly $(0,1)\in S[(1,0)]$, because $(1,1)\in S$. Now let $(a,b)\in T=R\times R$, thus there exist $x,y\in R$ such that $\sigma_{1}(x)=a$ and $\sigma_{2}(y)=b$, because $\sigma_{1},\sigma_{2}\in\operatorname{Aut}(R)$. Thus by definition of $S$ we conclude that $(\sigma_{1}(x),\sigma_{2}(x))=(a,\sigma_{2}(x))\in S$ and $(\sigma_{1}(y),\sigma_{2}(y))=(\sigma_{1}(y),b)\in S$. Therefore $(a,b)=(a,\sigma_{2}(x))(1,0)+(\sigma_{1}(y),b)(0,1)\in S[(1,0)]$. Thus $S[(1,0)]=T$. Now we prove that for each $(x,y)\in T\setminus S$, we have $S[(x,y)]=T$, which shows that $S$ is a maximal subring of $T$. Since $\sigma_{2}\in\operatorname{Aut}(R)$, there exists $c\in R$ such that $\sigma_{2}(c)=y$. This implies that $(\sigma_{1}(c),y)\in S\subseteq S[(x,y)]$ and a fortiori $(x-\sigma_{1}(c),0)\in S[(x,y)]$. Since $(x,y)\notin S$, we infer that $x_{0}:=x-\sigma_{1}(c)\neq 0$. Thus $Rx_{0}R=R$, because $R$ is a simple ring. Hence there exist $n\in\mathbb{N}$ and $r_{i},s_{i}\in R$, $1\leq i\leq n$, such that $\sum_{i=1}^{n}r_{i}x_{0}s_{i}=1$. From $\sigma_{1}\in\operatorname{Aut}(R)$, we deduce that there exist $a_{i},b_{i}\in R$, $1\leq i\leq n$, such that $\sigma_{1}(a_{i})=r_{i}$ and $\sigma_{1}(b_{i})=s_{i}$. It follows that
@@
This implies that $T=S[(1,0)]\subseteq S[(x,y)]$ and then $S[(x,y)]=T$. Consequently, $S$ is a maximal subring of $T$.
Conversely, assume that $S$ is a maximal subring of $T$. Suppose $I=R\times 0$ and $J=0\times R$. If either $I\subseteq S$ or $J\subseteq S$, then $S$ is of the forms which is given in $(1)$. Thus assume that $I$ and $J$ are not contained in $S$. Therefore $I_{1}=I\cap S=\{(x,o)\ |\ x\in S_{1}\}=S_{1}\times 0$ and $J_{1}=J\cap S=\{(0,y)\ |\ y\in S_{2}\}=0\times S_{2}$, where $S_{1}$ and $S_{2}$ are proper subsets of $R$. Clearly $I_{1}+J_{1}=S_{1}\times S_{2}\subseteq S$. Also note that since $S$ is a maximal subring of $T$ which does not contain $I$ and $J$, we deduce that $S+I=T=S+J$ and thus $S/I_{1}\cong R\cong S/J_{1}$. Consequently, $I_{1}$ and $J_{1}$ are maximal ideals of $S$. Now we have two cases: $(\bf a)$ If $(1,0)\in S$ (which is equivalent to $(0,1)\in S$), then $(1,1)=(1,0)+(0,1)\in I_{1}+J_{1}$, which means $I_{1}+J_{1}=S$ and therefore $S=S_{1}\times S_{2}$. Thus $S_{1}$ and $S_{2}$ are proper subring of $R$. But obviously $S\subsetneq R\times S_{2}\neq T$, which contradicts to the maximality of $S$. $(\bf b)$ Hence we may suppose that $(1,0),(1,0)\notin S$. It implies that $I_{1}+J_{1}=S_{1}\times S_{2}$ is a proper ideal of $S$ which contains $I_{1}$ and $J_{1}$. But as we see earlier, $I_{1}$ and $J_{1}$ are maximal ideals of $S$. Therefore we obtain that $I_{1}=J_{1}$. So $I_{1}=J_{1}=0$ and then $S\cong R$. Let $\sigma:R\longrightarrow S$ be a ring isomorphism, and for each $x\in R$, let $\sigma(x)=(\sigma_{1}(x),\sigma_{2}(x))$. Then clearly $\sigma_{1}$ and $\sigma_{2}$ are ring endomorphisms of $R$. Since $R$ is a simple ring we deduce that $\sigma_{1}$ and $\sigma_{2}$ are one-one. Finally note that from $S=\{(\sigma_{1}(x),\sigma_{2}(x))\ |\ x\in R\}$ and $(1,0)\notin S$, we have $S[(1,0)]=T$. It follows that $T=S+S(1,0)$ (note, $(1,0)$ is a central idempotent of $T$). To show that $\sigma_{1}$ and $\sigma_{2}$ are onto, let $x\in R$, then $(x,0)\in S+S(1,0)$. Consequently there exist $a,b\in R$ such that $(x,0)=(\sigma_{1}(a),\sigma_{2}(a))+(\sigma_{1}(b),\sigma_{2}(b))(1,0)$, which shows that $\sigma_{1}(b)=x$, i.e., $\sigma_{1}$ is onto. Analogously, $\sigma_{2}$ is onto, and hence we are done. ∎
We now present a main result of this paper concerning simple rings.
@@
Let $T$ be a simple ring which is not a division ring. Then $T$ has only finitely many maximal subring if and only if $T$ is finite (i.e., $T\cong\mathbb{M}_{n}(F)$, where $F$ is a finite field and $n\geq 2$)
@@
We first show that $\operatorname{{Max}_{\ell}}(T)$ is finite. Otherwise, let $M_{1},M_{2},\ldots$ be distinct maximal left ideals of $T$. Then for each $i$, by Theorem 2.7(1)(1), $A_{i}:=\mathbb{I}_{T}(M_{i})$ is a maximal subring of $T$ which contains $M_{i}$ (note, for each $i$, $M_{i}$ is not an ideal of $T$, because $T$ is simple ring which is not a division ring). Since whenever $i\neq j$, we have $M_{i}+M_{j}=T$, then we have $A_{i}\neq A_{j}$. Thus $T$ has infinitely many maximal subrings which is absurd. Thus $\operatorname{{Max}_{\ell}}(T)=\{M_{1},\ldots,M_{n}\}$. Now from the fact that $J(T)=0$, we deduce that $M_{1}\cap\cdots\cap M_{n}=0$. Consequently, $T$ embeds in $T/M_{1}\times\cdots\times T/M_{n}$. A fortiori, $T$ is a left Artinian ring. It follows that $T\cong\mathbb{M}_{n}(D)$, for some division ring $D$ and $n\geq 2$ (note, $T$ is not a division ring). If $D$ is an infinite division ring, then $\mathbb{M}_{n}(D)$, where $n\geq 2$, has infinitely many maximal left (right) ideals, by Corollary 2.4. Thus $D$ is finite and then $D$ is a finite field. This implies that $T$ is a finite ring, which complete the proof. ∎
We recall that for a field $K$, the ring $K\times K$ has only finitely many maximal subrings if and only if $K$ is finite, see [10, Corollary 3.5]. The following result generalizes this fact.
@@
Let $R$ be a simple ring. Then $R\times R$ has only finitely many maximal subrings if and only if $R$ is finite.
@@
It is trivial that if $R$ is finite then $R\times R$ has only finitely many maximal subrings. Conversely, assume that $R\times R$ has only finitely many maximal subrings. Theorem 3.4 implies that $R$ has only finitely many maximal subrings. If $R$ is not a division ring, then we are done by the previous theorem. Thus suppose that $R$ is a division ring and $R\times R$ has only finitely many maximal subrings. By the preceding comment, we may assume that $R$ is not a field, i.e., $C_{R}(R)\subsetneq R$. Now, let $R$ be an infinite ring. Then we have two cases: $(\bf a)$ If there exists a non-central element $x$ of $R$ which is algebraic over the center of $R$, then by Proposition 3.1, $R$ has infinitely many maximal subrings, which is impossible. $(\bf b)$ Hence assume that each non-central element $x$ of $R$ is not algebraic over the center of $R$. Let $x\in R$ be a non-central element and for each $n\geq 1$ and $a\in D$, let $\sigma_{n}(a)=x^{n}ax^{-n}$ be the inner automorphism of $R$ induced from $x^{n}$. If $\sigma_{n}=\sigma_{m}$ for some $n>m$, then $\sigma_{1}^{n-m}=1$ and therefore $x^{n-m}$ is a central element of $R$. This implies that $x$ is integral over the center of $R$, which contradicts to the assumption that $x$ is not algebraic over the center. Thus $\sigma_{n}\neq\sigma_{m}$ for each $n\neq m$. Consequently, by Theorem 3.4(2)(2), for each $n\geq 1$, the set $S_{n}=\{(x,\sigma_{n}(x))\ |\ x\in R\}$ is a maximal subring of $R\times R$. Obviously, whenever $n\neq m$, we have $S_{n}\neq S_{m}$. It follows that $R\times R$ has infinitely many maximal subrings, which is absurd. Thus $R$ is a finite, as desired. ∎
@@
By [12, Theorem 6.5.11], $D$ has at least $2^{\aleph_{0}}$ automorphisms. Hence by Theorem 3.4 for each $\sigma\in\operatorname{Aut}(D)$, $S_{\sigma}:=\{(x,\sigma(x))\ |\ x\in D\}$ is a maximal subring of $D\times D$. It is clear that $S_{\sigma}\neq S_{\sigma^{\prime}}$, whenever $\sigma$ and $\sigma^{\prime}$ are two distinct automorphisms of $D$. Hence $D\times D$ has at least $2^{\aleph_{0}}$ maximal subrings. ∎
@@
It is clear that if $(1)$ holds then $T$ has infinitely many maximal subrings. Now assume that $(2)$ holds, thus $T\cong R_{1}\times R_{1}$ and therefore by Proposition 3.6, we conclude that $T$ has infinitely many maximal subrings. Conversely, assume that $T$ has infinitely many maximal subrings, we prove that $(1)$ or $(2)$ holds. Suppose that $(2)$ does not hold, i.e., $R_{1}$ and $R_{2}$ are not isomorphic as rings. Thus $R_{1}$ and $R_{2}$ are homomorphically non-isomorphic rings. Consequently, by Lemma 2.9, we infer that each maximal subrings of $T$ is of the form $S_{1}\times R_{2}$ or $R_{1}\times S_{2}$, where $S_{i}$ is a maximal subring of $R_{i}$. Since $T$ has infinitely many maximal subrings, we obtain that either $R_{1}$ or $R_{2}$ has infinitely many maximal subrings and hence we are done. ∎
For a ring $T$, let us denote the sets $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$ by $\operatorname{{Max}_{\ell}}^{\prime}(T)$, similarly $\operatorname{{Max}_{r}}^{\prime}(T):=\operatorname{{Max}_{r}}(T)\setminus\operatorname{Max}(T)$ and $\operatorname{Max}^{\prime}(T):=\operatorname{Max}(T)\setminus\operatorname{{Max}_{\ell}}(T)=\operatorname{Max}(T)\setminus\operatorname{{Max}_{r}}(T)$. Note that if $P$ is a prime ideal of a ring $T$ and $I$ be a left (resp. right) ideal and $J$ is an ideal of $T$ such that $P=I\cap J$, then either $P=I$ or $P=J$. To see this, note that $JI\subseteq I\cap J\subseteq P$ (resp. $IJ\subseteq I\cap J\subseteq P$), therefore $I\subseteq P$ or $J\subseteq P$. Hence $P=I$ or $P=J$.
@@
    $\operatorname{{Max}_{\ell}}^{\prime}(T)$ and $\operatorname{{Max}_{r}}^{\prime}(T)$ are finite sets and in fact $|\operatorname{{Max}_{\ell}}^{\prime}(T)|\leq|\operatorname{RgMax}(T)|$ and $|\operatorname{{Max}_{r}}^{\prime}(T)|\leq|\operatorname{RgMax}(T)|$.